OpenStudy (anonymous):

PLEASE HELP ME!! Express each ratio as a unit rate. Round to the nearest tenth, if necessary. 11.) 343.8 miles on 9 gallons

jimthompson5910 (jim_thompson5910):

all you need to do is divide both parts by 9 343.8/9 = ??? 9/9 = 1

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

you have to convert miles to gallons I thought

jimthompson5910 (jim_thompson5910):

nope, just find the unit rate and you do this by dividing both parts by the second value

jimthompson5910 (jim_thompson5910):

besides, converting miles to gallons doesn't make sense since one is a distance and the other is a volume

OpenStudy (anonymous):

Ok one second.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

382

OpenStudy (anonymous):

38.2

jimthompson5910 (jim_thompson5910):

good, so 343.8 miles on 9 gallons turns into 38.2 miles on 1 gallon

jimthompson5910 (jim_thompson5910):

which means the unit rate is 38.2 miles per gallon

OpenStudy (anonymous):

thank you! can you help me with some more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

13) $52 for 8 tickets.

jimthompson5910 (jim_thompson5910):

so how would you do this use the last problem as a template to get you going if you're stuck

OpenStudy (anonymous):

6.5 ?

OpenStudy (anonymous):

$6.5 over 1 ticket

jimthompson5910 (jim_thompson5910):

so saying $52 for 8 tickets. is the same as saying $6.50 for 1 ticket (assuming all tickets are the same price)

jimthompson5910 (jim_thompson5910):

the unit rate is 6.5 dollars per ticket

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

next one?

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

450 miles in 8 hours

jimthompson5910 (jim_thompson5910):

divide both parts by 8 to get ....?

OpenStudy (anonymous):

56.25 per ticket

OpenStudy (anonymous):

i mean mi an hour

jimthompson5910 (jim_thompson5910):

good, 56.25 miles in one hour = 56.25 miles per hour

OpenStudy (anonymous):

thanks! I'll do the next 2 by myself and then can you help me with the next part?

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

19.) Which costs more per notebook, a 4-pack of notebooks for $3.98 or a 5-pack of notebooks for $4.99? explain

jimthompson5910 (jim_thompson5910):

find the unit rate for each so for the 4-pack of notebooks for $3.98 , the unit rate is... 3.98/4 = 0.995 which rounds to 1.00 so the unit rate for "4-pack of notebooks for $3.98" is $1 a notebook (pretty much)

jimthompson5910 (jim_thompson5910):

find the unit rate for "5-pack of notebooks for $4.99"

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

what does 0.998 round up to

jimthompson5910 (jim_thompson5910):

1.00

OpenStudy (anonymous):

its the same then

jimthompson5910 (jim_thompson5910):

you round to the nearest penny...ie, the nearest hundredth

jimthompson5910 (jim_thompson5910):

yep, both are the same deal

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

help me with the next ones

jimthompson5910 (jim_thompson5910):

if you consider fractions of a penny, they are different but you don't, which forces you to round landing you on the same numbers

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yep!

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

9.) Write a proportion that could be used to solve for each variable. Then solve 6 plums at $1 10 plums at d

jimthompson5910 (jim_thompson5910):

6 plums at $1 ---> 6/1 10 plums at d dollars ---> 10/d the two ratios are equal, so 6/1 = 10/d

jimthompson5910 (jim_thompson5910):

solve 6/1 = 10/d for d

OpenStudy (anonymous):

d=1.7?

jimthompson5910 (jim_thompson5910):

more like d = 1.667 which rounds to the nearest penny to get d =1.67

jimthompson5910 (jim_thompson5910):

so 10 plums cost $1.67

OpenStudy (anonymous):

11.) 3 packages at $53.67 7 packages at m

jimthompson5910 (jim_thompson5910):

3/53.67 = 7/m solve for m

OpenStudy (anonymous):

m= 125.23?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so 7 packages cost $125.23

OpenStudy (anonymous):

ok one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

On a set of architectual drawings for a school, the scale is 1/2 inch = 4 feet. Find the actual length of each room. 1.) Room: Classroom Drawing distance: 5 inches

jimthompson5910 (jim_thompson5910):

1/2 inch = 4 ft 0.5 inch = 4 ft

jimthompson5910 (jim_thompson5910):

5/x = 0.5/4 solve for x

OpenStudy (anonymous):

x=12 ft?

jimthompson5910 (jim_thompson5910):

5/x = 0.5/4 5*4 = 0.5x 20 = 0.5x 0.5x = 20 x = 20/0.5 x = 40

jimthompson5910 (jim_thompson5910):

so a drawing distance of 5 inches corresponds to a real life distance of 40 feet

OpenStudy (anonymous):

I understand I just wrote it wrong. I get how to do the problem but I don't get how you set it up.

OpenStudy (anonymous):

is it ft?

jimthompson5910 (jim_thompson5910):

yes it's 40 ft

jimthompson5910 (jim_thompson5910):

well 5 inches corresponds to x feet in real life so 5/x is one ratio 0.5 inches corresponds to 4 feet in real life, so the other ratio is 0.5/4

jimthompson5910 (jim_thompson5910):

the two ratios are equal, so that gives you 5/x = 0.5/4

jimthompson5910 (jim_thompson5910):

does the set up make more sense now?

OpenStudy (anonymous):

yes :)

jimthompson5910 (jim_thompson5910):

ok awesome

OpenStudy (anonymous):

one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

I can do the rest. Thank you so much :)

jimthompson5910 (jim_thompson5910):

you're welcome